# Understanding Differences in News Article Interaction Patterns on Facebook: Public vs. Private Sharing with Varying Bias and Reliability

### v2

@@ 4.2.2 Distribution of Public Interaction Shares
Key observation: Users exhibit higher deep interaction levels with content from the Most-unreliable class.
6 Correlation-based Analysis
Given the significant differences observed in interaction rates and deep interaction levels between public and private posts, we aim to determine if these variations stem from specific articles within each bias/reliability class that receive more or fewer private interactions or if certain articles consistently exhibit higher (deep) interaction rates, regardless of public or private sharing. To investigate this, we next examine the correlations between public post statistics and their combined counterparts.
Public vs. combined interactions: Table 3 summarizes the correlation statistics (between combined and public interactions) observed for each class using two different correlation metrics: Spearman and Kendall. Given that the normality assumption required for the Pearson correlation is not satisfied for the different classes, we opted for non-parametric correlation metrics. These metrics are better suited to capture monotonic relationships, account for outliers, and take into consideration the ranks of the data points.
Table 3: Correlation between public and combined interactions. (Bold/italics indicate highest/lowest correlation.)
Class Spearman Kendall
Bias Left 0.82 0.66
Center 0.74 0.59
Right 0.74 0.58
Reliability Most-unreliable 0.73 0.58
Unreliable 0.79 0.63
Reliable 0.76 0.61
Most-reliable 0.77 0.62
While different trends are observed for reliability classes, a consistent pattern emerges for the bias classes across both correlation metrics. The Left class exhibits the strongest correlations, followed by the Center class, and finally, the Right class with the weakest correlations. This suggests that articles with a left-leaning bias tend to have more similar sharing and interaction patterns in both public and private contexts, indicating higher similarity in engagement across different articles. In contrast, articles with a right-leaning bias show more variability in their public and private interactions, which may reflect a more diverse range of opinions or reactions to the content among readers. These observations are consistent with our observations in Section 4.2.2, which showed that the Right class saw much more diverse public interaction shares depending on the level of combined interactions of the article (see our bucket-based analysis).
Table 4: Correlations for deep interaction shares: public vs combined. (Bold/italics indicate highest/lowest correlation.)
Class Spearman Kendall
Bias Left 0.64 0.48
Center 0.62 0.47
Right 0.51 0.39
Reliability Most-unreliable 0.55 0.42
Unreliable 0.61 0.47
Reliable 0.64 0.49
Most-reliable 0.57 0.42
Deep interaction rates: We next consider the correlation between the deep interactions ratio for combined (DIRC) and public (DIRP) posts. Table 4 summarizes these results.
First, when comparing the correlations for deep interaction ratios across the bias classes, we observe that the Left and Center bias classes show higher correlations across both metrics, while the Right class exhibits weaker relationships. This pattern suggests that deep interactions, or more in-depth engagements with content, are more consistent between public and private contexts for left-leaning and centrist articles. Second, in terms of reliability, the correlations are relatively similar across classes, with the Reliable class showing the strongest relationships and the Most-unreliable class having the weakest. These findings indicate that the consistency of deep interactions across public and private contexts is less affected by articles’ perceived reliability than their political bias.
7 Discussion
Our research provides a comprehensive exploration of news interactions on Facebook, with a specific focus on discerning the nuances between public and combined interactions across diverse news classes. We developed a robust comparison methodology, crucial for future studies in this area. Our findings, presented in several key facets, provide valuable insights along several dimensions (discussed next).
Reliability and Bias: Several discoveries emerged when examining the reliability of news articles. Surprisingly, articles classified as Reliable garnered lower levels of engagement compared to those classified as Most-reliable or even various Unreliable categories. Furthermore, our investigation into the impact of biases unveiled that biased news, regardless of whether it leans Left or Right, consistently received higher levels of user interaction than more neutral (Center) content.
Public vs. Private: One of the central aspects of our study concerned the comparisons of the interaction patterns within the public vs. private spheres. What became evident was that articles classified as Most-reliable or Most-unreliable, representing the extremities of the reliability spectrum, attracted significantly higher proportions of engagement within private spaces compared to their public counterparts. Intriguingly, no statistically significant differences were observed between these two extreme groups here. This observation holds substantial implications for content moderators, policymakers, and researchers, prompting us to ponder whether an exclusive focus on the public facet of Facebook provides a comprehensive understanding of user interactions, particularly concerning news content classified as Most-unreliable.
Interaction Depth: Our research extended beyond interaction frequency to explore the depth of interactions, revealing noteworthy variations between the public and private spheres. Notably, statistically significant differences in interaction depth were observed, reinforcing the importance of considering both spheres in research endeavors. For example, the extremities of the reliability spectrum (Most-reliable and Most-unreliable) exhibited a tendency to elicit deeper interactions from users within the public sphere. However, this narrative shifted when examining combined interactions, indicating that only the Most-unreliable class triggered statistically significant deeper engagement from Facebook users. This underscores the intricate interplay between reliability, bias, and interaction depth, urging researchers to explore these dynamics beyond the confines of public engagement.
Ethical Considerations: It is essential to acknowledge the ethical dimension of our research. While our findings offer invaluable insights, they also hold the potential for misuse. To mitigate this risk, we undertook several precautions, such as selecting a diverse set of articles and news sources and maintaining objectivity in our analysis. We adhered to relevant privacy regulations, ensuring that no personally identifiable information was collected or disclosed. Transparency is of utmost importance to us, and we will provide all codes used in this paper, along with the list of news articles, in a camera-ready version. Although we cannot publish raw engagement data due to CrowdTangle terms of service, we will furnish attachment status information for URLs and posts through a lookup table with hashed keys. This approach enables responsible utilization of our research findings while minimizing the likelihood of negative consequences.
Beneficiaries: Our findings hold value for a diverse range of beneficiaries. Researchers, journalists, and policymakers can gain insights into news engagement patterns, aiding in informed decision-making. Social media content moderators can benefit by understanding where to focus their efforts. Moreover, our insights have relevance for platforms and media organizations, guiding them in shaping content strategies for more effective audience engagement.
In conclusion, our research sheds light on complex engagement dynamics on social media and emphasizes the responsible use of these findings for informed decision-making.
8 Related Works
@@
This research most closely relates to studies exploring user engagement dynamics with various social media content. Our main contribution lies in the novel analysis of interaction disparities between public and private news article sharing on Facebook, particularly concerning articles with varying bias and reliability. Examining interactions depth provides another level of depth to our analysis and findings.
@@
There exist studies that partially touch upon some dimensions highlighted in this paper. However, as discussed in the introduction, the vast majority of these works, particularly pertaining to Facebook, predominantly concentrate on the platform’s public sphere, neglecting to provide a comparative analysis with the private or combined sphere. To contextualize our study within the context of the existing literature and to highlight the novelty of our contribution, we next describe the most closely related works.
@@
One line of this prior research investigates user engagement with different classes of news, particularly regarding bias and reliability. For example, Edelson et al. [7], while focusing on the public sphere, conducted a large-scale study on user engagement with 7.5 million posts from 2551 publishers (Facebook pages) across both bias and reliability dimensions. Their results show that individual posts from non-misinformation news outlets tend to attract lower median engagement than misinformation. Weld et al. [34] carried out a similar investigation on Reddit rather than on Facebook. When considering the reliability parameter, Vosoughi et al. [35] analyzed a dataset of approximately 126,000 stories tweeted by over 3 million people more than 4.5 million times, finding that false news stories reached more people than true news and diffused significantly farther, faster, deeper, and more broadly than the truth across various categories of information. Notably, during the critical period of the COVID-19 pandemic and by doing a cross-national study, Altay et al. [36] delineated the surge in online news consumption, with credible news outlets witnessing a significant boost. Samory et al. [37] characterized the social media news sphere through user co-sharing practices by focusing on 639 news sources, both credible and questionable, and divided them into 4 clusters, and characterizing them according to the audience that shares their articles on Twitter and how the stylometric features used by each cluster is successful in getting users engagement. Lamot et al. [38] examined how news headlines are remediated on Facebook and how this affects user engagement. Finally, Boukes et al. [39] compared user-content interactivity and audience diversity across news and satire, finding differences in online engagement between satire, regular news, and partisan news. Considering the bias dimension, on the other hand, Wischnewski et al. [40] discovered that users are more inclined to share hyperpartisan news articles that coincide with their own political views. The bias towards the right party in sharing and engagement with news is another thesis that has been studied by previous works on both Twitter [41] and Facebook [42].
Other dimensions of user engagement have also been considered. For example, Aldous et al. [8] investigated the effects of topic and emotional factors on posting across five social media platforms (Facebook, Instagram, Twitter, YouTube, and Reddit). They found that user engagement is strongly influenced by the content’s topic, with certain topics being more engaging on a particular platform. By controlling the effect of the author (considering articles by the same author) over a one-year period, Maier et al. [43], among other things, found that highly emotional topics get a higher level of Facebook engagement. A large body of research has also focused on studying the role of sad and anger in social media news sharing and engagement (e.g., [44, 45]).
Finally, we note that profiling of the relation between user engagement dynamics and different content itself has paved the way for new applications. Examples include techniques and models that rely (partly) on user engagement dynamics (as features) for the detection of different content types. Such applications may benefit social media content moderators [46, 47]. Others have profiled the temporal dynamics of user engagement with different classes of news [12, 48].
@@
In summary, while previous research has partly explored some dimensions of user engagement with news articles, our study is the first to directly compare the dynamics of public and private sharing of news articles on Facebook, specifically examining the interactions and depth of engagement with articles of different levels of bias and reliability.
9 Conclusions
In this paper, we presented the first investigation of the differences in interaction patterns between public and private sharing of news articles on Facebook, focusing on articles with varying bias and reliability, as well as the depth of interactions. By analyzing a dataset of over 19K news articles (with more than 225M interactions on Facebook), employing the CrowdTangle browser addon for data collection, and carefully processing this data, we provided a comprehensive analysis of user interaction with news content on Facebook.
Our findings reveal significant differences in interaction patterns between public and private sharing across different classes of news. Notably, users tend to engage more deeply in private discussions compared to public ones, irrespective of the bias or reliability class of the news. We also found that Facebook users engage more deeply with the least reliable news and the right-biased news, both classes exhibiting higher deep interaction ratio levels. While this tendency may suggest that the content from these classes may be more polarizing or controversial, leading to deeper discussions and debates, further research (e.g., by interviewing users) is needed to determine the underlying factors driving this difference in engagement.
Our research findings offer valuable insights with significant implications for various stakeholders (e.g., researchers, journalists, policymakers, and social media content moderators) and deepen our understanding of how public and private spheres influence the dissemination and engagement with news articles on social media, enabling more informed decision-making. For example, our identification of content classes that generate deeper user engagement can provide actionable guidance to Facebook moderators, enabling them to focus their efforts effectively. Furthermore, the disparities we uncovered between public and private interaction patterns can guide future research that considers both spheres, promoting a more comprehensive approach. To facilitate such analysis, we present a robust methodology developed through active collaboration with CrowdTangle, which ensures a fair comparison between public and private spheres. This methodology equips researchers and content moderators with a valuable tool to navigate the evolving landscape of online news engagement.
In conclusion, by examining the differences in public and private sharing of news articles, this study offers valuable insights into the user engagement patterns associated with contents belonging to different bias and reliability classes, as well as highlights the importance of understanding the dynamics of news sharing and interaction both in the public and private spheres of social media platforms such as Facebook. While challenges remain, our methodology and discoveries offer valuable data-driven guidance toward addressing both spheres.
10 Statements and Declarations
10.1 Conflict of interest
The authors have no competing interests.
10.2 Author contributions
Both authors have made substantial contributions to the conceptualization, design, data collection, analysis, and writing of the research.
10.3 Acknowledgment
The authors thank the CrowdTangle team for their active feedback and help during this research.
10.4 Data and code availability
All the codes associated with this research will be made available with the camera-ready version. At the same time, to facilitate replicating the results and conducting similar research, the list of all news articles considered in this study and also the attachment status of all related posts will be published.

### v3

@@ 4.2.2. Distribution of Public Interaction Shares
Takeaway: Users exhibit higher deep interaction levels with content from the Most-unreliable class.
6. Topic-Based Analysis
In this section, we extend our investigation by conducting a topic-based analysis of news articles to assess whether engagement patterns vary across different content themes. For this analysis, we first extracted the textual content from the dataset news articles, a process that succeeded for 14,996 items. We then employed the ChatGPT 4o-mini model to classify each article into one of 14 predefined topics: Politics, Health & Medicine, Crime, Business  & Finance, Environment &  Climate, Entertainment, Education, Sports, Science, Lifestyle & Leisure, Religion, Technology, Arts &  Culture, and Food. This specific set of topics was selected to capture the diverse range of subject matters prevalent in contemporary news consumption, while also reflecting societal interests and trends.
Notably, the ordering of topics here mirrors the distribution observed in our dataset—ranging from the most frequent (Politics, with 7,956 articles) to the least frequent (Food, with 57 articles). However, it is important to note that these frequencies are influenced by two key factors: (1) the initial article selection criteria of Ad Fontes Media, and (2) the timeframe of our dataset, which spans a period heavily impacted by the COVID-19 pandemic, leading to an elevated presence of Health & Medicine content (the second most prevalent topic). Finally note that for articles where none of these predefined topics were clearly identifiable, the model was instructed to leave them uncategorized. In total, 14,650 articles (97.7%) were successfully classified into one of the 14 categories.
First, to assess whether our earlier findings hold within specific topics, we conducted a stratified analysis across topic groups. Our results indicate that the previously observed patterns remain, with statistical significance for several of the classes with enough samples. For example, in the case of the Politics topic the insights reported in both Sects. 4 and 5 remain statistically robust when considering the reliability dimension. As another example, for Health & Medicine (the 2nd largest topic), all the conclusions in Sects. 4 and 5 for both the bias and reliability dimensions are still statistically significant. These confirm the general validity of our broader findings within many of these topics.
Second, to explore how engagement varies across topics, we analyzed the distributions of the public interaction ratio for each topic. This analysis was conducted independent of the articles’ bias and reliability, as some topics had too few samples to allow for a breakdown by those categories. Figure 9 shows the distribution of the public interaction ratios (as a boxplot) for each of the topics, sorted based on their medians. From this figure, we can first see that Entertainment-related articles exhibit the highest public interaction ratio, indicating that content from this category is more likely to be engaged with in public discussions. This suggests that entertainment news, which often includes celebrity updates and viral stories, is inherently more shareable and widely disseminated in public spaces. On the other hand, Technology-related articles exhibit the lowest public interaction share, meaning they receive a disproportionately high share of engagement in private spaces. This may be attributed to the technical nature of these articles, which are often discussed within niche communities or shared in professional or interest-based private groups.
Figure 9. Boxplots of the public interaction ratio for different topics (sorted by median).
Figure 10. Heatmap of Kruskal-Wallis test results for topics’ public interaction ratios.
It is important to note that not all observed differences in public interaction ratios between topics in Figure 9 are statistically significant. To validate the statistical significance of these observed differences, we conducted pairwise Kruskal-Wallis tests between topic groups. Figure 10 presents a heatmap illustrating the statistical significance of differences in public interaction ratios across topics. Orange-colored cells here indicate significant differences ($p<0.1$), while gray cells indicate non-significant comparisons.
This analysis confirms that Entertainment-related content has a significantly higher public interaction share compared to all other topics, whereas Technology-related content exhibits significantly lower public engagement compared to most categories. Conversely, topics such as Food, we could not find statistically significant differences in public interaction ratios compared to other topics (except for Entertainment). It is worth noting that we selected a p-value threshold of 0.1 for our analysis, as stricter thresholds (e.g., 0.01) did not yield many statistically significant differences between topics. This more lenient threshold was chosen primarily to compensate for the limited sample size for certain topics in our dataset.
7. Top Publishers Analysis
To further contextualize our findings, we analyze the public interaction ratio (PIR) and the deep interaction ratio (DIRC) for the top publishers in our dataset. In this context, “top publishers” refers to those with the highest number of articles in our (and Ad Fontes Media) dataset. For this analysis, we excluded Yahoo, as it primarily functions as a news aggregator and does not produce original content.
We begin by examining the public interaction ratio across the top publishers. Figure 11(a) presents a boxplot of the public interaction ratio for each outlet, sorted by median values. For easier interpretation, the color of each “box” indicates the bias class (i.e., Left, Center, or Right) associated with each publisher (based on the labeling of the articles they have published). It is worth noting that some outlets, such as Politico and NBC News, may shift between the Left and Center classifications over time due to editorial changes or evolving media landscapes. From this figure, we observe that New York Post exhibits the lowest public interaction ratio, suggesting that a larger proportion of its engagement occurs in private spaces. This could be attributed to the nature of its audience, which may prefer engaging with its content in more private settings, such as closed groups. Additionally, the sensationalist reporting of the New York Post might encourage discussions that users feel more comfortable having in less-visible online spaces.
Next, we investigate which of these publishers generate the deepest user engagement, as measured by the deep interaction ratio (DIRC). Figure 11(b) illustrates the DIRC distributions for the top publishers, again sorted by median values. Notably, New York Post and Fox News, both from the Right class, exhibit the highest median DIRC values, followed by CNN and New York Times from the Left class. At the lower end of the spectrum, we find NPR and Reuters, both from the Center class. In general, we observe that publishers with the highest DIRC values tend to be biased ones, while those from the Center class have the lowest values. The probability of this pattern occurring if we randomly sort the 10 publishers is approximately 2.22%. This observation may suggest that users are more likely to engage deeply with content from publishers that have a clear political leaning, compared to those that maintain a more neutral stance.
(a) Boxplots of the public interaction ratios (sorted by median).
(b) Boxplots of the deep interaction ratios (sorted by median).
Figure 11. Comparison of (a) public interaction ratio and (b) deep interaction ratio for the top-10 outlets.
8. Related Works
@@
This research most closely relates to studies exploring user engagement dynamics with various social media content. Our main contribution lies in the novel analysis of interaction disparities between public and private news article sharing on Facebook, particularly concerning articles with varying bias and reliability. Examining interactions depth adds another layer to our analysis and findings.
@@
There exist studies that partially address some dimensions highlighted in this paper. However, as discussed in the introduction, the vast majority of these works, particularly pertaining to Facebook, predominantly concentrate on the platform’s public sphere. To contextualize our study within the context of the existing literature and to highlight the novelty of our contribution, we next describe the most closely related works.
@@
One line of this prior research has studied the differences in the engagement levels of different bias or reliability (or both) classes on different platforms including Facebook (Edelson et al., 2021; Hiaeshutter-Rice and Weeks, 2021; Altay et al., 2022), Reddit (Weld et al., 2021; Bond and Garrett, 2023), and Twitter(Mohammadinodooshan and Carlsson, 2024; Vosoughi et al., 2018). For example, Edelson et al. (Edelson et al., 2021), while focusing on the public sphere, conducted a large-scale study on user engagement with 7.5 million posts from 2,551 publishers (Facebook pages) across both bias and reliability dimensions. Their results show that individual posts from non-misinformation news outlets tend to attract lower median engagement than misinformation, aligning with our findings. Weld et al. (Weld et al., 2021) carried out a similar investigation on Reddit rather than on Facebook. When considering the reliability parameter, Vosoughi et al. (Vosoughi et al., 2018) analyzed a dataset of approximately 126,000 stories tweeted by over 3 million people more than 4.5 million times, finding that false news stories reached more people than true news and diffused significantly farther, faster, deeper, and more broadly than the truth across various categories of information. Notably, during the critical period of the COVID-19 pandemic and by doing a cross-national study, Altay et al. (Altay et al., 2022) delineated the surge in online news consumption, with credible news outlets witnessing a significant boost. Samory et al. (Samory et al., 2020) characterized the social media news sphere through user co-sharing practices by focusing on 639 news sources, both credible and questionable, and divided them into 4 clusters, and characterizing them according to the audience that shares their articles on Twitter and how the stylometric features used by each cluster is successful in getting users engagement. Lamot et al. (Lamot et al., 2022) examined how news headlines are remediated on Facebook and how this affects user engagement. Finally, Boukes et al. (Boukes et al., 2022) compared user-content interactivity and audience diversity across news and satire, finding differences in online engagement between satire, regular news, and partisan news. Considering the bias dimension, on the other hand, Wischnewski et al. (Wischnewski et al., 2021) discovered that users are more inclined to share hyperpartisan news articles that coincide with their own political views. The bias towards the right party in sharing and engagement with news is another thesis that has been studied by previous works on both Twitter (González-Bailón et al., 2022) and Facebook (González-Bailón et al., 2023).
Research has also examined other dimensions of user engagement, with, for example, Aldous et al. (Aldous et al., 2019) showing that content topics influence engagement. Studies have found that emotional content generates higher Facebook engagement, with Maier et al. (Maier, 2015) demonstrating this effect while controlling for author influence. Finally, some works have considered the time factor and analyzed temporal patterns of engagement with different news classes (Vassio et al., 2022).
@@
In summary, while previous research has partly explored some dimensions of user engagement with news, our study is the first to directly compare the dynamics of public and private sharing of news articles on Facebook.
9. Conclusions
In conclusion, this study makes several important contributions to the understanding of news engagement on social media. By developing a robust methodology through collaboration with the CrowdTangle team, we present the first comprehensive comparison of engagement dynamics between Facebook’s publicly tracked content and overall platform interactions. Our analysis of over 19K news articles reveals that public engagement patterns often fail to reflect platform-wide behavior, with significant implications for research methodology and content moderation strategies.
Our findings challenge several assumptions about news engagement on Facebook. First, users consistently engage more deeply with content in less public spaces, regardless of the content’s bias or reliability classification. Second, particular attention should be paid to right-biased and least reliable news content, which generate notably deeper engagement across both public and less public spheres. Third, we highlight some important limitations of focusing solely on Facebook’s public sphere. For example, the extremities of the reliability spectrum (Most-reliable and Most-unreliable) exhibited a tendency to elicit deeper interactions from users within the public sphere. However, this narrative shifted when examining combined interactions, indicating that only the Most-unreliable class triggered statistically significant deeper engagement from Facebook users. This finding suggests that research conclusions drawn solely from public data may not accurately represent overall platform engagement patterns, as not only the magnitude of engagement differs between public and private but also the relative ordering of how different content types engage users.
These insights carry significant implications for multiple stakeholders. For researchers, our findings underscore the importance of considering engagement disparities between public and less public spheres when studying social media behavior, particularly for news content at different ends of the bias and reliability spectrums. For content moderators, our results suggest the need to reevaluate strategies based solely on publicly visible interactions. For policymakers, our findings highlight the importance of considering both public and less public spheres when developing policies for social media governance.
Future research should investigate the underlying factors driving these engagement patterns, particularly through user interviews, behavioral studies, and analysis of post comments. While challenges remain in understanding social media dynamics, our findings illuminate significant differences in how users interact with news across Facebook’s diverse spheres, highlighting the complexity of news-sharing behavior on social media platforms.
Acknowledgements.
The authors express their gratitude to CrowdTangle for providing the Facebook data. They also extend their thanks to the anonymous reviewers for their insightful comments that helped improve the paper. This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.